The age of a father is 10 less than the sum of the ages of his two sons. 7 years ago his age was 3 times the sum of the ages of his sons find their present ages
Found 2 solutions by fractalier, ikleyn:
Answer by fractalier(6550)

f = s - 10
f-7 = 3(s-7)
s-10 - 7 = 3s - 21
s - 17 = 3s - 21
s = 2
Impossible. The father cannot be -8 years old.
You have posted the wrong question as it is written. Please check it.

Answer by ikleyn(52798)

.
The age of a father is 10 less than the sum of the ages of his two sons. 7 years ago his age was 3 times
the sum of the ages of his sons find their present ages
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

f = s - 10,       (1)
f-7 = 3*(s-14).   (2)     <<<<--- "14" must be distracted, not "7" !

--->  (s-10) -7 = 3s - 42

s - 17 = 3s - 42,

2s = 42 - 17 = 25.


s = 12.5



Hello, the author! 


1. Change the data in your problem to allow "s" to be integer.


2. From the given data it is impossible to determine the age of the two sons. Too few data.